When 0.149 mol of carbon is burned in a closed vessel with

8.73 g

of oxygen, how many grams of carbon dioxide can form?
Answer #1

The reaction is:

C (s) + O2 (g) = CO2 (g)

Moles of C = 0.149 mole

Moles of O2 = Mass/ Molar mass = 8.73 g / (32 g / mole) = 0.273 mole

The reaction between C and O2 occurs at 1:1 mole ratio. So, here as the moles of C is less than the moles of O2, C is the limiting reagent.

Mole ratio of C : CO2 = 1:1

Moles of CO2 formed = Moles of C reacted = 0.149 mol

Mass of CO2 formed = 0.149 mole * Molar mass = 0.149 mole * 44 g/ mole = 6.56 g